Question 2

The primary advantage of the related samples design is that it controls for ________ that may artificially increase the observed differences between means in other types of studies.

Group of answer choices

degrees of freedom

MD

individual differences

the effect of the treatment

Question 3

A manager would like to know if the performance of sales representatives significantly (α = .05) improves over the first 6 months of employment. For 5 new sales representatives, the number of sales in the first month is recorded and the number of sales in the sixth month is recorded. What is the null hypothesis for this repeated measures t-test?

Group of answer choices

µD = 0

µD > 0

µ1 = µ2

µD ≤ 0

µ ≤ 5

µ1 > µ2

Question 7

A researcher obtains t = 2.80 for a repeated-measures study examining the difference in reading ability before and after completing a reading comprehension course among a sample of n = 16 participants. What is the value of r2?

Group of answer choices

r2 = .149

r2 = .329

r2 = .343

cannot be determined with the information given.

Question 8

For the repeated measures t-test, the value of r2 tells you ___.

Group of answer choices

the proportion of differences in the dependent variable which can most likely be attributed to the independent variable

the proportion of differences in the independent variable which can most likely be attributed to the dependent variable

the proportion of differences between two measures of the dependent variable which can be accounted for by individual differences

the proportion of differences between two measures of the independent variable which can be accounted for by individual differences
Answer #1

Question 2

Answer:

individual differences

----------

Question 3

Answer:

µD = 0

-------------

Question 7

t = 2.80

n = 16

df = n-1 =15

r² = t²/(t²+df) = 2.8²/(2.8²+15) = 0.343

Answer: r2 = 0.343

-------------

Question 8

Answer:

the proportion of differences in the independent variable which can most likely be attributed to the dependent variable